OpenStudy (anonymous):

Find the next term of the following sequence. 25, 10, 4, ... I got 0 am I right

OpenStudy (zehanz):

I think it is meant to be a geometric series. Every next term is 2.5 times smaller. That way, you don't arrive at 0...

OpenStudy (anonymous):

my choices are 0 -5 8/5

OpenStudy (zehanz):

Well, if you divide 4 by 2.5 (= 5/2) you can also multiply 4 by 2/5 and that gives you on of the options...

OpenStudy (anonymous):

so is it 8/5

OpenStudy (zehanz):

Yes!

OpenStudy (anonymous):

hey I need help I have to do the problem on here cuz idk how to do it on the post thing

OpenStudy (anonymous):

\[\sqrt[3]{-27n^2^7}\]

OpenStudy (zehanz):

Do you mean:\[\sqrt[3]{-27(n^2)^7}\]

OpenStudy (anonymous):

no n^(27)

OpenStudy (anonymous):